DFR Telecoms Diary


Wednesday, January 05, 2005

I went to the DFR and had a productive day. Bernard and I nearly have the 10+50 working. I also did a lot of recovery of cable round the museum and out near the booking hut. I took the shop decorations down and helped John Metherall get the booking office sign down for repainting.

Wednesday, January 12, 2005

I finished off the 10+50 which is now working well. I helped Norman reroute a power cable to a socket in the shop. I helped John M put up a couple of notices in the museum. I also sorted out the C&W group to weld me up a couple of brackets.

Wednesday, January 19, 2005

I moved the phone from beside the office door to above the exchange window, helped Bernard sort out the eyeball on the 10+50 and put PBX strapping on 878 and 879. We had a check of the PA levels on the phone connections but had to pack it in as we did not have sufficient time to complete the job. We need to set all the levels from scratch.

Wednesday, January 26, 2005

I fixed the telephone exchange PA phone and then spent the afternoon trying to get the levels right. They seemed OK when we left but it will be interesting if things are still Ok next week.

Wednesday, February 02, 2005

I spent the day putting up a phone and the wall clock in the museum. The clock was made to work though.

Wednesday, February 09, 2005

I fitted the PA button and replaced the phone in the museum office and did a couple of faults. We also sorted out the trip for the exchange power so that I can reorganise the over ride switch later.

Wednesday, March 02, 2005

I fitted the new light in the "parcel's office" in the museum, adjusted the new clock and started on the shift of the mains cut off switch for the exchange. Bernard laid flooring in the exchange.

Wednesday, March 09, 2005

I went to the DFR where I fitted the mains isolation switch outside the exchange, fitted a voltmeter to the MDF and helped John Metherall to put the ticket office sign back up.

Wednesday, March 16, 2005

I spent the day somewhat inconclusively trying to shift the TV monitor leads to the new office and testing the battery on discharge. Two batteries are faulty. I fitted lots of labels around the work we have done lately.

Friday, March 18, 2005

We met Chris Hall of the MHR. We picked up quite a bit of gear at Medstead and then went back to Ropley station to collect some more. However we have to return to collect the rest of the gear. We had a gentle run home as the car was really weighted down.

Wednesday, March 23, 2005

We shifted the CCTV monitor to the new office and got it working. Bernard fitted the heater under rack A but it was too warm to need it. I recovered some 50pr in the loft and we sorted out the battery situation. I also fitted an extra mains socket in the exchange.

Wednesday, March 30, 2005

During the day I increased the clock current, changed the PA time clock and fitted a cut off relay to the 10+50 exchange lines. We are still having discussions about changing the battery and rectifier.

Saturday, April 02, 2005

We went to Alresford to meet Chris Hall of the MHR. We had tea with him and picked up some gear. The place was crowded as it was a Thomas the Tank day. We went on to Medstead to get the rest of the gear. Then it was a gentle and slow ride home in the slow lane.

Wednesday, April 06, 2005

We changed out the two faulty cells in the battery for the MHR cells, changed a fuse in the 10+50 power unit, and ran mains cables across the south end of the workshop ready for terminating later.

Thursday, April 21, 2005

The phone rang with John Hatton wanting me to check on the ringer at the DFR. When we got there all was OK.

Wednesday, April 27, 2005

I cleared a couple of faults in the exchange and started cabling to clear up the rack common services at Norchard.

Tuesday, May 03, 2005

Parkend Exchange under construction at my home

We had a meeting with Peter Adams and Tony Smith, the contractor, during the morning. It was agreed that we have to move to a system that gets electrical work certified. Afterwards I went back to the common services for rack A2, however I got sidetracked on the ring start condition finally finding out that the ringer starts if there is a condition on the time pulse start lead.

Wednesday, May 11, 2005

I proved a handset faulty on the cordless phone used for the PA. I also recovered some spare cable and blocks and an external bell in Bowley's Box. I attended a meeting with Mike Cornick and Derek Spinks in the signal box to discuss the telephone arrangements needed to make a line spare for the signal box.

Wednesday, May 18, 2005

I spent a fairly frustrating day on the rearrangements of the BT lines. I did however get the new signal box phone and bell connected back to the exchange.

Wednesday, May 25, 2005

I worked at the DFR sorting out the BT lines and changing them about a bit. Bernard helped by getting the lockable box in the kiosk and connecting up the phones there.

Wednesday, June 01, 2005

The DFR was an odd day. We checked Parkend for the position of the proposed BT telephone. I finished off getting the BT line to the signal box. I had to sort out Denise's phone and answering machine. We also threw some rubbish away from the caboosh.

Wednesday, June 08, 2005

I spent most of the day at Parkend dealing with the BT Installer. We got the line working by about four. I got the phone and extension bell fitted but the phone needs a new one really.

Wednesday, June 15, 2005

We spent the day on the Museum Door Alarm. Nearly finished but the second sounder would not work. A small mod needed next week.

Wednesday, June 22, 2005

I got the museum alarm finished and then John H and I checked the BT lines. We can't finish without getting hold of a PIN number. After lunch we walked the line towards Parkend to look at the duct. I moved it closer to the fence over most of its length. We finished by looking at display lighting in the museum.

Wednesday, June 29, 2005

I spent most of the day surveying the work necessary to get cable to Whitecroft. During the afternoon John Roff and I walked to Whitecroft and caught the DMU back.

Wednesday, July 06, 2005

I spent the day sorting the stores with Bob Hawker for phase 2 of the Parkend cable. I also got the barring done on the kiosk phone and also put metering on the four museum demo phones.

Wednesday, July 13, 2005

We spent the day shifting heavy items. We cleared a space in the Severn and Wye Room and then filled it with Christmas toys. Then we laid out a cable along the track to the end of the existing duct and then got half of it covered by duct and we also got a rope in the duct.

Wednesday, July 20, 2005

We spent the day continuing with the duct and cable laying. We are nearly complete for this phase but we still need cabinets and terminations.

Wednesday, July 27, 2005

I spent the day with Bernard finishing off the ducting, reconnecting ducting that had come apart, putting back up a SW notice and sorting out the route from the MDF to DP3a for a relief cable. At four, John Hatton and I met Steve Harris at the junction to sort out the pieces for the three cabinets we want at the moment. I dropped a steel bar on my ankle at the Junction, quite painful and swollen by the time I had got home.

Wednesday, August 03, 2005

I spent the day on the relief cable to DP3A with Bernard and for some time with Norman. We got the cable in and terminated on the MDF.

Wednesday, August 10, 2005

Overcast to start but the sun came back out again and it turned pretty warm. I went to the DFR picking up Les on the way. Bernard and I spent most of the morning tapping out the cable to DP3A and getting it terminated. Now we have pairs to most places except the Junction. Later we turned to erecting the posts and tiebars for the cabinet at DP5B, hard work as ended up digging down onto a concrete plinth or similar.

Wednesday, August 17, 2005

Bernard and I erected the posts and bars for the DP by the crossing and in the afternoon I cleared faults in the exchange.

Wednesday, August 24, 2005

I spent the day getting A2 common services changed over to separate links on the test panel. I picked up Jean on the way home before going to the PO.

Wednesday, August 31, 2005

Bernard and I with help from John H and Norman got the first two cabinets erected onto their posts. I also altered the PG wiring in the exchange to get the panel working correctly. I also terminated the rest of the D sides in DP5A and got the pairs jumpered through and tested. At least the first length looks to be fine.

Wednesday, September 14, 2005

I helped Bernard a bit with getting the DP5B cabinet ready to terminate and I mended the guards torch. I had two DMU rides up to Whitecroft surveying what is to come.

Wednesday, September 21, 2005

I got DP5B completed. I also put up the backboard in DP5C and walked along to sort out the exact position for DP5D at the far end of the existing duct.

Wednesday, September 28, 2005

I spent the day terminating DP5C. I couldn't finish as the rain came in and pushed us indoors.

Wednesday, October 05, 2005

Only Ivor and I were there. I finished the terminating and rebuilding the cabinet at DP5C. The cable though shows some low insulation, hope it's just the far end.

Wednesday, October 12, 2005

I cut the duct at DP5D and shortened the cable back to the DP position. I cleaned up the end and taped it up and finished just as the rain started. The insulation value is now OK. I spent the rest of the day sorting out the cabinet for DP5D and got to the point where I needed a door. The rain was hammering down by then so that was the end of work.

Wednesday, October 19, 2005

I dug a couple of post holes, shifted some duct along the line to the next required place, shopped at Travis Perkins, made up bolts etc for the DP5D cabinet and cut a board for the cabinet interior for signwriting etc. Bernard took over the cabinet refurbishment.

Wednesday, October 26, 2005

I spent the day not too profitably. We did get the cabinet and all the bits up to DP5D but it rained and we had to give up and return to Norchard. I did have quite a discussion about the Index of Appliances with John H inorder to try to find out what he wanted. We made the bolts up for fixing the tie bars to the concrete posts as well. I left just after three as the rain prevented any more work being undertaken.

Wednesday, November 02, 2005

I went to Parkend to meet Stuart Bearne and Peter Adams to discuss the Parkend Exchange. It was left that the exchange would be stored at Parkend until required and that it would probably go into a new building at the back of the goods shed. I am to provide a specification for the room. On the way back I checked that the cabinet 5D was actually up so I went on to collect the cover and took it back to fit it.

Wednesday, November 09, 2005

I went off to terminate the cable in DP5D and got that completed by lunch. After lunch I went to Parkend to look at the position for the telephone and to measure the room in the goods shed. Then I put a plug on a small strip light so that we could see what they are like in the museum cabinets. John Roff and I also moved three lengths of duct up to the crossing ready to be dug in.

Wednesday, November 16, 2005

Ivor announced that he was leaving due to a medical condition reducing his driving. John Hatton seems to have recovered some drive in getting things done. I spent the day on taking over Ivor's records, repairing the two electrical defects in the exchange, and putting lights in a display cabinet.

Wednesday, December 14, 2005

We had had a main power fuse blow over the weekend so John H was sorting that out. I got to change out a blue socket and plug on the platform that had welded itself together with heat. I also got to use the new tester from BT to find out the cable lengths we are dealing with. I also checked the line resistances at the same time. Now I can measure the cable lengths we have in store.

Wednesday, December 21, 2005

I spent the morning getting the lights to work on the DMU posts and the afternoon measuring the cable lengths we have.

Wednesday, December 28, 2005

I fixed a DMU light that had had the bulb taken, replaced a fuse in the 10+50 power unit and spent most of the day sorting out teles etc from the loft and caboosh.

Thursday, December 29, 2005

We spent most of the day sorting out the 300 type telephone bits from the DFR and making up complete teles. We did get two 332 types wiped up properly and have several more to go if I can find cords etc.

Return to the Diary Menu


Page provided by John Bathgate

This page was last updated on
9th August 2013